A disk takes 1 ms per track seek time , and the is head is currently on track 70. Assume it takes 5 ms to service a request on the head is at right track. The following set of requests are pending , given as (ID , track) in order of arrival. How much time does it take to service all the requests if the elevator algorithm is used and the current direction is down?

Requests ( A, 80) , (B,75) , (C,82) , (D,90) , (E,45)

Total time?
Answer #1

total time= (70-45) + (45-0) +(75-0) +(80-75) +(82-80) +(90-82) + (6*5) ; [6= number of tracks to be serviced, 5= service time]

=25 +45+ 75 + 5 +2 + 8 +30

=190 ms.